# Hyperparameter Transfer with Mixture-of-Experts Layers

### v2

@@ Abstract
Mixture-of-Experts (MoE) layers have emerged as an important tool in scaling up modern neural networks by decoupling total trainable parameters from activated parameters in the forward pass for each token. However, sparse MoEs add complexity to training due to (i) new trainable parameters (router weights) that, like all other parameter groups, require hyperparameter (HP) tuning; (ii) new architecture scale dimensions (number of and size of experts) that must be chosen and potentially taken large. To make HP selection cheap and reliable, we propose a new parameterization for transformer models with MoE layers when scaling model width, depth, number of experts, and expert (hidden) size. Our parameterization is justified by a novel dynamical mean-field theory (DMFT) analysis. When varying different model dimensions trained at a fixed token budget, we find empirically that our parameterization enables reliable HP transfer across models from 51M to over 2B total parameters. We further take HPs identified from sweeping small models on a short token horizon to train larger models on longer horizons and report performant model behaviors.
@@ 1 Introduction
This motivates the study of HP transfer (Yang and Hu, 2022; Yang et al., 2022; Bordelon and Pehlevan, 2022; Bordelon et al., 2023; Yang et al., 2023b; Bordelon and Pehlevan, 2025; Bordelon et al., 2024; Dey et al., 2025; Bergsma et al., 2025a; Wang and Aitchison, 2025), a suite of techniques for obtaining performant HPs in large models directly from good HPs found by tuning substantially smaller models. This HP extrapolation, or transfer, is determined through a parameterization, a set of theoretically-motivated rules that predict how variations in each scaling dimension of interest (e.g. model width, depth) should modify raw HP values.
@@
MoE parameterization.  We extend the CompleteP parameterization (Dey et al., 2025) (developed for dense transformers) to include MoE-specific scaling rules, allowing for scaling up not only depth and width but also the number and size of experts (Section˜3.3) at fixed active expert sparsity (Section˜3.2). Furthermore, we find empirically that training stability in MoEs is more sensitive to constant-scale HPs, multipliers that are treated as $\Theta(1)$ in our parameterization (see Apdx. D.1). In contrast, tuning these HPs in dense models can improve performance but does not seem necessary to ensure stability (Bordelon et al., 2023).
@@
Empirical validation of HP transfer. We systemically verify (over different datasets and MoE sparsity levels) that our parameterization enables reliable transfer on both the initialization scale (init. std. $\sigma$) and learning rate (LR $\eta$) when varying width, depth, expert count, and expert size (Figure˜2 and Figure˜4) on a fixed token budget (1B). We verify HP transfer by tuning base models with 38M activated parameters and scale up to (around) 2B total parameters (Figure˜2 and Figure˜14).
@@ 2 Related backgrounds
HP transfer and DMFT.  Traditional approaches to HP tuning at large scale require either grid-searching in a large model or extrapolating from power laws fit to a family of smaller models (Li et al., 2025; Zhou et al., 2026). Instead, our approach seeks the direct (rule-based) transfer of optimal HP from small models to larger ones. The core behind transfer is parameterizations that stabilize the forward and backward passes across scales. First studied for scaling the width of MLPs, this stability condition resulted in two types of parameterizations: the Neural Tangent parameterization (NTP) (Hayou et al., 2022; Jacot et al., 2020; Roberts et al., 2022) and the Mean-Field parameterization (MFP) (Mei et al., 2018; Chizat et al., 2019; Rotskoff and Vanden-Eijnden, 2022). Such stability conditions have also been studied via the modular norm (Large et al., 2024) and effective field theory (Dinan et al., 2023; Yaida, 2022; Roberts et al., 2022).
@@
Dynamical mean-field theory (DMFT) provides a way to theoretically analyze HP transfer by showing which parameterizations cause training dynamics of finite networks to converge to a well-defined limit. This framework traces the evolution (in infinite-sized networks) of mean-field model statistics, such as inner-product kernels for hidden activations and gradients. DMFT results were first derived for MLPs (Bordelon and Pehlevan, 2022) and later extended for multi-head self-attention (Bordelon et al., 2024), transformers in the infinite-depth limit (Bordelon et al., 2023), and finite-width correction (Bordelon and Pehlevan, 2025).
@@ 3.1 Model setup
where $g_{i}(h)\triangleq\sigma(r_{i})\in\mathbb{R}$ are (un-normalized) mixing coefficients, $\sigma$ is a non-linear (sigmoid) function, and
@@
with a hidden layer of size $\alpha_{\operatorname{ffn}}\cdot n_{\operatorname{embd}}\in\Omega(n_{\operatorname{embd}})$, and $W^{(i)}_{\operatorname{up}},W^{(i)}_{\operatorname{down}}\in\mathbb{R}^{n_{\operatorname{embd}}\times\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}$. To be close to our theoretical analysis (Section˜4), our setup disentangles mixing coefficients for different experts (following (Wang et al., 2024; ArceeAI et al., 2026)): for each expert $i$, router $g_{i}(h)$ only depends on $W_{\operatorname{router}}^{(i)}\in\mathbb{R}^{n_{\operatorname{embd}}}$. Practically, we did not observe significant differences in model performance across scale between different types of mixing coefficients (e.g., topK-then-softmax, softmax-then-topK).
@@ 3.2 Scaling that preserves sparsity versus topK
We have three practical motivations for this choice. First, (assuming perfect balancing) in a batch of $B$ tokens, each MoE expert sees only $\kappa B$ tokens, whereas attention and routing modules see the full $B$ tokens. This suggests HPs will not transfer across different $\kappa$ (this is partly confirmed in Figure 11). Second, for large deployments, it has been empirically reported that the range of optimal sparsity $\kappa$ is often hardware-controlled (StepFun and others, 2025) and bounded below. Finally, a constant $\kappa$ is natural in the mean-field analysis (Sec. 4) as it represents conditioning on a constant-probability event when integrating over the mean-field measure of experts. Our theory suggests that HPs will transfer at fixed $\kappa$ and our empirical results support this point.
@@
Table 1: Parameter groups and their scaling rules for the MoE module at each layer as per derivation in Section˜$3.$
@@ 3.3 Proposed MoE Parametrization
We defer the heuristic justification of $\eta_{\operatorname{bias}}$ to Appendix B. As before, max-update principles place no requirement on initialization so long as $b\in O(1)$. However, taking into account expert load balancing, we conveniently initialize biases at zero so no one expert disproportionately receives tokens at initialization
@@
Note that the dependence of $\sigma_{\operatorname{init}}(W_{\operatorname{down}}^{(i)})$ concerning $\alpha_{\operatorname{ffn}}$ is distinct from the standard fan-in initialization (see Figure˜11 for a simple ablation). A heuristic justification (see also (Chizat, 2025)) is to imagine $n_{\operatorname{embd}}=1$ then treat $\alpha_{\operatorname{ffn}}$ as the intermediate “width” in a two-layer MLP under the mean-field parametrization (as opposed to NTP, which yields the standard fan-in initialization).
@@ 4 Dynamical mean field theory
The existence of a well-defined mean-field limit in the training dynamics is a strong justification of HP transfer (Yang and Hu, 2022; Bordelon et al., 2023, 2024). To theoretically establish that our parameterization enables a scale-invariant limit (and consequently admits HP transfer across jointly scaling model dimensions), we outline here a novel approach to studying the training dynamics for a deep residual MoE model [^1] While we focus on analyzing a MoE-block only residual stream for the context of this work, we see no obstruction to deriving complete DMFT equations for the full sparse transformer architecture by combining with existing analysis of multi-head self-attention layers in (Bordelon et al., 2024). in which each input token $x\in\mathbb{R}^{D}$ is processed by first up-projection $h^{(0)}=W_{\mathrm{embed}}x\in\mathbb{R}^{n_{\mathrm{embd}}}$, then passing through $L$ residual MoE feedforward layers
$h^{(\ell+1)}=h^{(\ell)}+L^{-1}f_{\mathrm{MoE}}^{\ell}(h^{(\ell)})\in\mathbb{R}^{n_{\mathrm{embd}}},\quad\ell=1,\ldots,L$
@@ 5.3 Tradeoff between expert count versus expert size
6 Conclusion and future directions
@@
    As training dynamics in early stages (such as our experiments in 1B token scale) versus later stages (e.g., the “compute-optimal” horizon) can be very different, it would be interesting to expand the HP transfer to longer training horizons that are practical.
@@
    We’d also like to remark that it is not known what “compute-optimal” rules can be practically applied for MoEs (Clark et al., 2022), as (a) Mixture-of-Expert layers achieve significantly better performance compared to dense models under the same FLOP budget, so Chinchilla exponents may not apply, and (b) even when FLOP-matched, MoEs take up significantly more demanding hardware resources due to sparsity and routing, suggesting the necessity of a better way of evaluating compute.
    We take our present work on HP transfer as a first step to derive practical scaling laws for MoE models (see (Clark et al., 2022; Krajewski et al., 2024; Li et al., 2025; Zhao et al., 2025) for prior MoE laws in different settings).
@@
    Finally, we left out discussions of hyperparameters beyond learning rate and init standard deviation, such as batch size, Adam betas, and LR schedules, all of which are often interconnected and can have a more significant impact as data scale expands.
@@ Appendix A More experiment details
In this section we lay-out some further experiment details from Section˜5. We train decoder-only MoE models described in Section˜3.1. Our attention, embedding, and normalization setups are based on the public Nano-GPT repo (Karpathy, 2023) and standard GPT-2 style tokenizer (Radford et al., 2019) with a vocabulary size of 50304. For all optimization, we use standard Adam betas $\beta_{1},\beta_{2}=0.9,0.95$ and a negligible Adam $\varepsilon=10^{-12}$. We use a fixed $d_{\operatorname{head}}=64$ (while scaling $\operatorname{num\_head}\propto n_{\operatorname{embd}}$ for large embedding width) for multi-head self-attention following (Dey et al., 2025). We use $QK^{T}/d_{\operatorname{head}}$ (as opposed to standard $\sqrt{d_{\operatorname{head}}}$) for the normalization of self-attention following (Yang et al., 2022). For experiments with scaling models, our base model has dimensions of base $n_{\operatorname{embd}}=512$, base $\alpha_{\operatorname{ffn}}=1$, base $L=8$, and we up-scale relevant multipliers via the prescribed parametrization in Table˜1 (we point our parameterizations for all non-FFN HPs and LN HPs to Table 1 in (Dey et al., 2025)). We train our models using the standard autoregressive cross-entropy loss (i.e. the next token prediction objective) and always report the log perplexity score. As with the standard Nano-GPT (Karpathy, 2023) setup, we use pre-LayerNorm, tied embeddings, absolute learned position embeddings, and GELU nonlinearity. Because router matrices takes up $O(n_{\operatorname{embd}}^{-1})$ fraction of total FFN parameter counts, we may ignore them and derive total parameters as
@@
Our experiments are run on H100s and are bit-wise deterministic. Our code is available at https://github.com/Petyrrrrr/MuP_MOE/tree/complete.
Appendix B Heuristic LR derivation for expert biases
We derive a simple heuristic on why bias learning rate should not be adjusted when scaling the total number of experts while fixing sparsity. Consider the following pseudo-code, where we assume that the pre-activated $\sigma^{-1}(g_{i})$’s are i.i.d. Gaussian, (existing) biases are uniformly distributed in a fixed range, and updated biases are uniformly distributed in a fixed range. This represents the training phase where load imbalance is $\Theta(1)$ proportion of total tokens. The goal is that, regardless of the total number of experts, for a fixed set of gradient distribution, a constant bias learning rate enables constant expert activation set change. See comments below for how we set-up the heuristics.
@@ Appendix C Weight decay
Appendix D Other experiment results
@@ D.1 Constant-scale HP tuning for the base model
In (Yang and Hu, 2022; Yang et al., 2022), the authors pointed out that good HP transfer is crucially based on the “close to optimality” of the scaling. Without specific care, HPs found from tuning small models can be “contaminated” by finite-width effects, which fail to yield useful transfer (Figure 17, (Ghosh et al., 2025)). This motivates tuning constant-scaled multipliers on relevant HPs, whose benefits for dense models were remarked in (Everett et al., 2024; Mlodozeniec et al., 2025; Ghosh et al., 2025) that not only promote better loss (in the base model) but likely contribute to more reliable transfer. In our experiments with MoE layers, we found that balancing constant-scaled hyperparameters not only leads to better performance, but it is specifically crucial for stable training dynamics (across different random seeds) on large learning rates. In practice, we enable two constant multipliers for each MoE parameter groups for learning rate and initialization. While each group of weights has a three relevant HPs: learning rate, initialization, and multiplier, in the so-called abc-parameterization (Yang and Hu, 2022), only two degrees of freedom are present. We will take all weight multipliers to be one, so only initializations and learning rates are at play. We report loss results on ablation in Figure˜9 and load balancing ablation in Figure˜9.

### v3

@@ Abstract
Mixture-of-Experts (MoE) layers have emerged as an important tool in scaling up modern neural networks by decoupling total trainable parameters from activated parameters in the forward pass for each token. However, sparse MoEs add complexity to training due to (i) new trainable parameters (router weights) that, like all other parameter groups, require hyperparameter (HP) tuning; (ii) new architecture scale dimensions (number of and size of experts) that must be chosen and potentially taken large. To make HP selection cheap and reliable, we propose a new parameterization for transformer models with MoE layers when scaling model width, depth, number of experts, and expert (hidden) size. Our parameterization is justified by a novel dynamical mean-field theory (DMFT) analysis. When varying different model dimensions trained at a fixed token budget, we find empirically that our parameterization enables reliable HP transfer across models from 51M to 2B total parameters. We further take HPs identified from sweeping small models on a short token horizon to train larger models on longer horizons and report performant model behaviors.
@@ 1 Introduction
This motivates the study of HP transfer (Yang and Hu, 2022; Yang et al., 2022; Bordelon and Pehlevan, 2022; Bordelon et al., 2023; Yang et al., 2023b; Bordelon and Pehlevan, 2025; Bordelon et al., 2024; Dey et al., 2025; Bergsma et al., 2025a; Wang and Aitchison, 2025; Wu et al., 2026), a suite of techniques for obtaining performant HPs in large models directly from good HPs found by tuning substantially smaller models. This HP extrapolation, or transfer, is determined through a parameterization, a set of theoretically-motivated rules that predict how changing each scaling dimension of interest (e.g. model width, depth) should modify raw HP values.
@@
MoE parameterization.  We extend the CompleteP parameterization (Dey et al., 2025) (developed for dense transformers) to include MoE-specific scaling rules, allowing for scaling up not only depth and width but also the number and size of experts at fixed active expert sparsity (Section˜3.3 and Section˜3.2) without re-tuning HPs at each scale.
@@
Empirical validation of HP transfer. We systematically verify (over different datasets and MoE sparsity levels) that our parameterization enables reliable transfer on both the initialization scale (standard deviation $\sigma$) and learning rate (LR $\eta$) when varying width, depth, expert count, and expert size (Figure˜2 and Figure˜4) on a fixed token budget (1B). We verify HP transfer by tuning base models with 38M activated parameters and scale up to (around) 2B total parameters (Figure˜2 and Figure˜14). During training, we report that stability in MoE pre-training is more sensitive to constant-scale HPs, multipliers that are treated as $\Theta(1)$ in our parameterization (Apdx. D.1). In contrast, balancing constant-scale HPs in dense models can improve performance but does not seem necessary to ensure stability (Bordelon et al., 2023).
@@ 2 Related backgrounds
HP transfer.  Traditional approaches to HP tuning at scale require either grid-searching in a large model or extrapolating from power laws fit to a family of smaller models (Li et al., 2025; Zhou et al., 2026). Instead, our approach seeks the direct (rule-based) transfer of optimal HP from small models to larger ones. The core behind transfer is parameterizations that stabilize the forward and backward passes across scales. First studied for scaling the width of MLPs, this stability condition resulted in two types of parameterizations: the Neural Tangent parameterization (NTP) (Hayou et al., 2022; Jacot et al., 2020; Roberts et al., 2022) and the Mean-Field parameterization (MFP) (Mei et al., 2018; Chizat et al., 2019; Rotskoff and Vanden-Eijnden, 2022). Such stability conditions have also been studied via the modular norm (Large et al., 2024) and effective field theory (Dinan et al., 2023; Yaida, 2022; Roberts et al., 2022).
@@
DMFT Dynamical mean-field theory (DMFT) provides a way to theoretically analyze HP transfer by showing training dynamics of finite networks to converge to a well-defined limit. This framework traces the evolution (in infinite-sized networks) of mean-field model statistics, such as inner-product kernels for hidden activations and gradients. Earlier DMFT analysis were derived for MLPs (Bordelon and Pehlevan, 2022), Deep ResNets (Bordelon et al., 2023), multi-head self-attention (Bordelon et al., 2024) and transformers in the infinite-depth limit (Bordelon et al., 2023).
@@
[…]
After the first version of this work, we noticed independent and concurrent work (Vankadara et al., 2026), which (also) studies the DMFT limit of MoE models under a fixed sparity. The main difference between our studies is that the notion of a scale-invariant limit in (Vankadara et al., 2026) is somewhat different from ours, and hence the necessity of a more sophisticated analysis in their regimes.
@@ 3.1 Model setup
where $g_{i}(h)\triangleq\sigma(r_{i})\in\mathbb{R}$ are (un-normalized) mixing coefficients, where $\sigma$ is a non-linear (sigmoid) function, and
@@
with a hidden layer of size $\alpha_{\operatorname{ffn}}\cdot n_{\operatorname{embd}}\in\Omega(n_{\operatorname{embd}})$, and $W^{(i)}_{\operatorname{up}},W^{(i)}_{\operatorname{down}}\in\mathbb{R}^{n_{\operatorname{embd}}\times\alpha_{\operatorname{ffn}}n_{\operatorname{embd}}}$. Following (Wang et al., 2024; Dai et al., 2024; Singh et al., 2026) as well as a line of open-source models using similar architectures, our setup disentangles weights of expert mixing: for each expert $i$, $g_{i}(h)$ only depends on $W_{\operatorname{router}}^{(i)}\in\mathbb{R}^{n_{\operatorname{embd}}}$. Practically, we did not observe significant differences in model performance across scale between different types of mixing coefficients (e.g. softmax weighting, see Figure˜18).
@@ 3.2 Scaling that preserves sparsity versus topK
We have three practical motivations for this choice. First, (assuming perfect balancing) in a batch of $B$ tokens, each MoE expert sees only $\kappa B$ tokens, whereas self-attention and routing see the full $B$ tokens. This suggest HPs will not transfer across different $\kappa$ (partly confirmed in Figure 11). Second, for large deployments, it has been empirically reported that the practical range of $\kappa$ is often hardware-bottlenecked by communication cost (Liu et al., 2024; StepFun and others, 2025) and thus bounded below. Finally, a constant $\kappa$ is natural in the mean-field analysis (Sec. 4) as it represents conditioning on a constant-probability event when integrating over the mean-field measure of experts. Our theory suggests that HPs will transfer at fixed $\kappa$ and our empirical results support this point. For applications with a bottlenecked $n_{\operatorname{act}}$, our recipe is still helpful via (a) start with a fixed target sparsity and scale up granularity; or (b) fix an expert configuration and scale up width/depth; or (c) apply distillation after training a model with more active parameters.
Table 1: Parameter groups and their scaling rules for the MoE module at each layer as per derivation in Section˜$3.$ See also Table˜2 in the Appendix for a complete version.
@@ 3.3 Proposed MoE Parametrization
We defer the justification of $\eta_{\operatorname{bias}}$ to Appendix B. As before, max-update principles place no requirement on initialization so long as $b\in O(1)$. However, taking into account expert load balancing, we conveniently initialize biases at zero so no one expert disproportionately receives tokens at init.
@@
Note that the dependence of $\sigma_{\operatorname{init}}(W_{\operatorname{down}}^{(i)})$ concerning $\alpha_{\operatorname{ffn}}$ is distinct from the standard fan-in initialization (see Figure˜11 for a simple ablation). A heuristic justification (see also (Chizat, 2025)) is to treat $n_{\operatorname{embd}}=1$ and $\alpha_{\operatorname{ffn}}$ as the intermediate “width” in a standard two-layer MLP under the mean-field parametrization (as opposed to NTP, which yields the standard fan-in initialization).
@@ 4 Dynamical mean field theory
The existence of a well-defined mean-field limit in the training dynamics is a strong justification of HP transfer (Yang and Hu, 2022; Bordelon et al., 2023, 2024). To theoretically establish that our parameterization enables a scale-invariant limit (and consequently admits HP transfer across jointly scaling model dimensions), we outline here a novel approach to studying the training dynamics for a deep residual MoE model [^1] While we focus on analyzing a MoE-block only residual stream for the context of this work, we see no technical obstruction to deriving complete DMFT equations for the full sparse transformer architecture by combining with existing analysis of multi-head self-attention layers in (Bordelon et al., 2024). in which each input token $x\in\mathbb{R}^{D}$ is processed by first up-projection $h^{(0)}=W_{\mathrm{embed}}x\in\mathbb{R}^{n_{\mathrm{embd}}}$, then passing through $L$ residual MoE feedforward layers
$h^{(\ell+1)}=h^{(\ell)}+L^{\!-1}f_{\mathrm{MoE}}^{\ell}(h^{(\ell)})\in\mathbb{R}^{n_{\mathrm{embd}}},\ell=0,\dots L-1$
@@ 5.3 Tradeoff between expert count versus expert size
6 Conclusions and future directions
@@
    As training dynamics in early stages (such as our experiments in 1B token scale) versus later stages (e.g., the “compute-optimal” horizon) can be very different, it would be interesting to expand the HP transfer to longer training horizons. We also left out discussions of hyperparameters beyond learning rate and init standard deviation, such as batch size, Adam betas, and LR schedules, all of which are interconnected and can have a stronger impact as training horizon expands.
@@
    While we include finite-dimensional analysis for our DMFT results in Appendix˜E, full non-asymptotic DMFT analysis (or HP transfer) is still largely open (see also (Bordelon and Pehlevan, 2023, 2025)). We verify that our theory works practically (e.g. that base $n_{\operatorname{embd}}=512$ suffices for HP transfer towards the infinite width limit) through experiments and leave the theoretical study of small-width transfer to future work (see also (Ghosh et al., 2025)).
    Furthermore, while our techniques in the DMFT analysis extend to SGD or SignGD naturally, a rigorous DMFT treatment on advanced optimizers such as Adam or Muon remain largely open.
@@
    Finally, we remark that it is not known what “compute-optimal” rules can be practically applied for MoEs (Clark et al., 2022), as (a) Mixture-of-Expert layers achieve significantly better performance compared to dense models under the same FLOP budget, so Chinchilla exponents may not apply, and (b) even when FLOP-matched, MoEs take up significantly more demanding hardware resources due to sparsity, suggesting the necessity of a better way of evaluating compute.
    We take our present work on HP transfer as a first step to derive practical yet rigorous scaling laws for MoE models (see (Clark et al., 2022; Krajewski et al., 2024; Li et al., 2025; Zhao et al., 2025) for prior MoE laws in different settings).
Impact Statement
@@ Appendix A More experiment details
In this section we lay-out some further experiment details from Section˜5. We train decoder-only MoE models described in Section˜3.1. Our attention, embedding, and normalization setups are based on the public Nano-GPT repo (Karpathy, 2023) and standard GPT-2 style tokenizer (Radford et al., 2019) with a vocabulary size of 50304. For all optimization, we use standard Adam betas $\beta_{1},\beta_{2}=0.9,0.95$ and a negligible Adam $\varepsilon=10^{-12}$. We use a fixed $d_{\operatorname{head}}=64$ (while scaling $\operatorname{num\_head}\propto n_{\operatorname{embd}}$ for large embedding width) for multi-head self-attention following (Dey et al., 2025). We use $QK^{T}/d_{\operatorname{head}}$ (as opposed to standard $\sqrt{d_{\operatorname{head}}}$) for the normalization of self-attention following (Yang et al., 2022). For experiments with scaling models, our base model has dimensions of base $n_{\operatorname{embd}}=512$, base $\alpha_{\operatorname{ffn}}=1$, base $L=8$, and we up-scale relevant multipliers via the prescribed parametrization in Table˜1 (we point our parameterizations for all non-FFN HPs and LN HPs to Table 1 in (Dey et al., 2025)). We train our models using the standard autoregressive cross-entropy loss (i.e. the next token prediction objective) and always report the log perplexity score. As with the standard Nano-GPT (Karpathy, 2023) setup, we use pre-LayerNorm, tied embeddings, absolute learned position embeddings, and GELU nonlinearity. Because router matrices takes up $O(n_{\operatorname{embd}}^{-1})$ fraction of total FFN parameter counts, we may ignore them and count the total parameters as
@@
Our experiments are run on H100s and are bit-wise deterministic. The entire set of experiments in our paper took around 5000 GPU-Hrs for H100s (the precise number of GPU-Hrs greatly depends on GPU-communication hardware). See the supplementary materials for empirical implementation details.
Appendix B Learning rate of expert biases
We derive a simple argument on why bias learning rate should not be adjusted when scaling the total number of experts while fixing sparsity. Consider the following pseudo-code, where we assume that the pre-activated $\sigma^{-1}(g_{i})$’s are i.i.d. Gaussian, (existing) biases are uniformly distributed in a fixed range, and updated biases are uniformly distributed in a fixed range. This represents the training phase where load imbalance is $\Theta(1)$ proportion of total tokens. The goal is that, regardless of the total number of experts, for a fixed set of gradient distribution, a constant bias learning rate enables constant expert activation set change. See comments below for how we set-up the heuristics.
@@ Appendix C Weight decay
Appendix D Other design details and experiment results
@@ D.1 Constant-scale HP tuning for the base model
Making constant-scale tuning in different types of weight matrices is a standard and studied practice for stability and performance (Everett et al., 2024; Mlodozeniec et al., 2025). Even in dense models, when training is inherently more stable, performance-driven implementations still make efforts to tuning and report benefits. For MoEs, we find that a minimal level of tuning is not only better for performance but also necessary for stability. In terms of HP transfer, it is also more reasonable to scale up from a base model (on which tuning is cheap and efficient) that is both stable and performant at a small scale, as base model misalignments/suboptimality could likely get exaggerated when scaling up (even if the scaling recipe is correct).
In (Yang and Hu, 2022; Yang et al., 2022; Ghosh et al., 2025), the authors pointed out that good HP transfer is crucially based on the “close to optimality” of the scaling. Without specific care, HPs found from tuning small models can be “contaminated” by finite-width effects, which fail to yield useful transfer (Figure 17, (Ghosh et al., 2025)). This motivates tuning constant-scaled multipliers on relevant HPs, whose benefits for dense models were remarked in (Everett et al., 2024; Mlodozeniec et al., 2025; Ghosh et al., 2025) that not only promote better loss (in the base model) but likely contribute to more reliable transfer. In our experiments with MoE layers, we found that balancing constant-scaled hyperparameters not only leads to better performance, but it is specifically crucial for stable training dynamics (across different random seeds) on large learning rates. In practice, we enable two constant multipliers for each MoE parameter group for learning rate and initialization. While each group of weights has a three relevant HPs: learning rate, initialization, and multiplier, in the so-called abc-parameterization (Yang and Hu, 2022), only two degrees of freedom are present. We will take all weight multipliers to be one, so only initializations and learning rates are at play. We report loss results on ablation in Figure˜9 and load balancing ablation in Figure˜9.